A group of students estimated the length of one minute without reference to a watch or​ clock, and the times​ (seconds) are listed below. Use a 0.01 significance level to test the claim that these times are from a population with a mean equal to 60 seconds. Does it appear that students are reasonably good at estimating one​ minute?

75 90 48 71 49 30 71 74 74 60 73 79 99 95 77

Determine test statistic

Determine P Value

State the final conclusion that addresses the original claim
Answer #1

Hypothesis:
H0 : mu = 60
Ha: mu not equasl to 60

test statistics:

xbar = 71 , sigma = 18.2718, n =15

t = (x - mu)/(sigma/sqrt(n))
= ( 71 - 60)/(18.2718/sqrt(15))
= 2.3316

p value = 0.0352

P(t > 2.3316) = 1 - P(x< 2.3316)
= 1 - 0.9824
= 0.0176

p value = 2 *0.0176 = 0.0352


fail to reject H0 as p value > 0.01

It Does not appear that students are reasonably good at estimating one​ minute